OpenStudy (dtan5457):

Create a polynomial function of at least degree 3, with rational coefficients and roots of:

OpenStudy (dtan5457):

\[x1=\frac{ 3-\sqrt{2} }{ 2 }\]\[x2=\frac{3+\sqrt{2} }{ 2 }\]\[x3=2i\]

OpenStudy (dtan5457):

@DanJS @Nnesha @Directrix

Nnesha (nnesha):

i have 3 minutes O_O

OpenStudy (dtan5457):

but you can take five

OpenStudy (dtan5457):

@Zale101

Nnesha (nnesha):

do you have to do this today

Nnesha (nnesha):

tag some people i gtg sorry and i know its gonna take at least 10 mintues

OpenStudy (alekos):

Just create 3 terms out of the roots and multiply it out

OpenStudy (dtan5457):

How do I do that?

OpenStudy (dtan5457):

@jim_thompson5910

OpenStudy (dtan5457):

@Nnesha got time now?

OpenStudy (alekos):

still there dtan?

Nnesha (nnesha):

we will do this today

OpenStudy (alekos):

I can go thru this now

OpenStudy (dtan5457):

go ahead

OpenStudy (dtan5457):

@alekos

OpenStudy (alekos):

OK lets call the roots a,b and c

OpenStudy (dtan5457):

alright

OpenStudy (alekos):

so we have three terms (x-a), (x-b) and (x-c) which when we multiply together we will get our cubic polynomial

OpenStudy (alekos):

(x-a)(x-b)(x-c) = dx^3 + ex^2 + fx + g

OpenStudy (dtan5457):

so i do multiply them all out in fraction form?

OpenStudy (alekos):

we just need to find the rational coefficents d,e,f and g

OpenStudy (dtan5457):

so what should a and b be?

OpenStudy (alekos):

lets start with (x-a)(x-b) where a = (3-sqrt2)/2 and b = (3+sqrt2)/2

OpenStudy (alekos):

x^2 - 3x + (11/4 + 3/sqrt2)

OpenStudy (dtan5457):

where did you get the 11/4

OpenStudy (dtan5457):

?

OpenStudy (alekos):

(3-sqrt2)/2 x (3+sqrt2)/2 =1/4(9 - 3sqrt2 + 3sqrt2 -2) = 1/4 x 7 =7/4 my first answer was incorrect

OpenStudy (alekos):

so we have x^2 - 3x +7/4

OpenStudy (alekos):

follow so far?

OpenStudy (dtan5457):

i see what your doing, where does the -2 come from though? =1/4(9 - 3sqrt2 + 3sqrt2 -2)<<

OpenStudy (dtan5457):

oh wait nevermind got it

OpenStudy (alekos):

-sqrt2 x sqrt2 = -2

OpenStudy (dtan5457):

square root of 4 ))

OpenStudy (dtan5457):

should that be 2i?

OpenStudy (alekos):

no

OpenStudy (dtan5457):

ok

OpenStudy (dtan5457):

carry on

OpenStudy (alekos):

if we now multiply (x^2 - 3x +7/4) with (x -2i) we get a cubic, but not all the co-efficients are real. so there must be another root probably -2i

OpenStudy (dtan5457):

so still there is 3 roots?

OpenStudy (alekos):

can't be . if you have 2i as a root then there must be a conjugate root which would be -2i

OpenStudy (dtan5457):

alright i get a better idea of this now. but it's late, ill look at this AGAIN tomorrow. thanks

OpenStudy (alekos):

ok. just multiply (x+2i)(x-2i) and then multiply the result with our quadratic above

OpenStudy (alekos):

you'll get a polynomial with degree 4 and with real co-efficients